# Bayesian optimisation of poloidal field coil positions in tokamaks

### v1

@@ 5 Discussion and outlook
To increase the applicability and extend this BO framework to ongoing and future PF coil design projects, a number of avenues of future work can be considered. For example, incorporating additional objectives and constraints should be a trivial task and could be used to help find PF coil sets that further improve divertor performance (e.g. maximising flux expansion) or help vertically stabilise the core plasma. It could also be of interest to include the PF coil thicknesses (in the input space) so that designers could extract even more financial cost savings by reducing the total cross-sectional area of the PF coils. This is something the current framework could be easily adapted to support. Beyond the current framework, analysis could also be carried out to investigate whether the coil sets found for the flat-top phase of the plasma pulse remain optimal during the ramp-up and ramp-down phases.
@@ Acknowledgements
The authors would like to thank Agnieszka Hudoba for providing the baseline STEP data files.
This work was funded by the EPSRC Energy Programme [grant number EP/W006839/1]. To obtain further information, please contact . For the purpose of open access, the author(s) has applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising.
@@
Data availability
The code scripts and data used in this paper will be made available in due course.
Declarations
The authors have no conflicts of interest to declare.
@@ Appendix A Coil location normalisation
An illustration of a PF coil and its permissible zone are shown in Figure A.1.
Figure A.1: Illustration of a rectangular PF (dark grey), with its centroid and half-width/height marked, and its permissible zone (red), with corner vertices marked.

### v2

@@ V Discussion and outlook
To increase the applicability and extend this BO framework to ongoing and future PF coil design projects, a number of avenues of future work can be considered. For example, incorporating additional objectives and constraints should be a trivial task and could be used to help find coilsets that further improve performance. For example, one could try to maximise flux expansion to improve divertor performance or include PF coil shaping (in the input space) to try to extract further financial cost savings. While this would increase the dimensionality of the BO problem, the current framework can be readily adapted to support this via dimensionality reduction techniques (Malu, Dasarathy, and Spanias, 2021; Constantine, Dow, and Wang, 2014; Grosnit et al., 2021).
Data Availability Statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.
@@ Acknowledgements
The authors would like to thank Agnieszka Hudoba for providing the baseline STEP data files and Theodore Brown for discussions around BO.
This work was funded by the EPSRC Energy Programme [grant number EP/W006839/1]. To obtain further information, please contact PublicationsManager@ukaea.uk. For the purpose of open access, the author(s) has applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising.
@@ Appendix A Coil location normalisation
An illustration of a PF coil and its permissible zone are shown in Figure 10.
Figure 10: Illustration of a rectangular PF (dark grey), with its centroid and half-width/height marked, and its permissible zone (red), with corner vertices marked.